OpenStudy (anonymous):

Find the 9th term of the arithmetic sequence. a10 = 100 and a20 = 50; a9 = _____.

hartnn (hartnn):

know the n'th term formula ?

OpenStudy (anonymous):

no, i just know find the differece and go with it. I work it out by hand it takes forever. But in this case I can't do that

hartnn (hartnn):

For Arithmetic sequences, \(\huge a_n =a_1 +(n-1)d \) Note it down, a1 = 1st term d = common difference n = number of terms

OpenStudy (anonymous):

but I don't know the common distance in this problem... That's why I posted it lol

hartnn (hartnn):

a10 = 100 100 = a1 + 9d got this ?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

a20 = 50 can you make another equation ?

OpenStudy (anonymous):

yeah 50=a1+19d

hartnn (hartnn):

100 = a1 + 9d 50 = a1 +19d 2 equations, 2 unknowns Subtract them,and, you will get the d

OpenStudy (anonymous):

How do I subtract them if they arent equal...

hartnn (hartnn):

you never subtracted 2 equations ? 100-50 = a1- a1 + 9d -19d

hartnn (hartnn):

simplify that

hartnn (hartnn):

tried ?

OpenStudy (anonymous):

50=-10d??

OpenStudy (anonymous):

d=-5?

hartnn (hartnn):

correct! now use any one of the equation to get a1

OpenStudy (anonymous):

so my equation would be 100=a1-5?? a1=95??

hartnn (hartnn):

100 = a1+9d 100 = a1 - 45 ....

OpenStudy (anonymous):

a1=55

hartnn (hartnn):

no...

OpenStudy (anonymous):

145...

hartnn (hartnn):

100 = a1-45 a1 = 100+45 yes

hartnn (hartnn):

now we need, a9, so we plug in n=9 in \(\huge a_n =a_1 +(n-1)d\)

hartnn (hartnn):

and you have already calculated a1 and d

OpenStudy (anonymous):

105

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

yay thank you!!!

hartnn (hartnn):

welcome ^_^